# Robust methods for Darcy–Forchheimer equationsDas, Hutridurga, Pani, Ruiz-Baier

### v1

@@ 4.1 Robust discrete solvability
First, we prove the discrete inf-sup condition, which is an essential ingredient of the proof of the parameter-robust unique solvability Theorem, discussed in this subsection. We adapt the same framework as in the continuous case to the present scenario.
@@
For the existence of ${\mathcal{S}_{d}}\in{\mathcal{L}}\bigl((\mathrm{W}_{h,\Gamma_{p}}^{1,q}(\Omega))^{\prime},{\mathbf{L}}^{r}_{h}(\Omega)\bigr)$ with $q\in\{2,\;\frac{3}{2}\}$, let $\phi_{h}\in\mathrm{W}^{1,r}_{h,\Gamma_{p}}(\Omega)$ for a given $g_{h}\in(\mathrm{W}_{h,\Gamma_{p}}^{1,q}(\Omega))^{\prime}$ with $\frac{1}{q}+\frac{1}{r}=1$ be a unique solution of the following weak formulation
@@
owing to the Banach–Nečas–Babuška theorem for elliptic problems in Banach spaces [13]. Then, it suffices to define $\mathcal{S}_{d}(g_{h}):=\nabla_{h}\phi_{h}$, where $\nabla_{h}\phi_{h}\in\mathbf{L}_{h}^{r}(\Omega)$. The continuous dependence on data coming from (4.7) also implies that $\|\mathcal{S}_{d}(g_{h})\|_{0,r;\Omega}\lesssim\|g_{h}\|$. This, in turn, shows $\mathcal{S}_{d}\in{\mathcal{L}}\bigl((\mathrm{W}^{1,q}_{h,\Gamma_{p}}(\Omega))^{\prime},{\mathbf{L}}_{h}^{r}(\Omega)\bigr)$, for $q=2$ and $q=\frac{3}{2}$ corresponding to $r=2$ and $3,$ respectively.
@@
$\|q_{h}\|_{{\widehat{\mathrm{Q}}(\mathcal{T}_{h})}}\leq{\beta_{d}}\sup_{\boldsymbol{v}_{h}\in\mathbf{V}_{h}}\frac{(q_{h},\mathop{\mathrm{div}}\nolimits\boldsymbol{v}_{h})}{\|\boldsymbol{v}_{h}\|_{\mathbf{V}}}.$
@@
$\sup_{\boldsymbol{0}\neq\boldsymbol{v}_{h}\in\mathbf{V}_{h}}\!\!\frac{({\mathop{\mathrm{div}}\nolimits}\,\boldsymbol{v}_{h},p_{h})}{\|\boldsymbol{v}_{h}\|_{\mathbf{V}}}$ $\geq\sup_{g_{h}\in\mathrm{Q}_{h}^{\prime}}\frac{({\mathop{\mathrm{div}}\nolimits}[\mathcal{S}_{d}(g_{h})],p_{h})}{\|\mathcal{S}_{d}(g_{h})\|_{\mathbf{V}}}=\sup_{g_{h}\in\mathrm{Q}_{h}^{\prime}}\frac{\langle g_{h},p_{h}\rangle}{\|\mathcal{S}_{d}(g_{h})\|_{\mathbf{V}}}$
$\geq\|\mathcal{S}_{d}\|^{-1}_{\mathcal{L}(\mathrm{Q}_{h}^{\prime},\mathbf{V}_{h})}\sup_{g_{h}\in\mathrm{Q}_{h}^{\prime}}\frac{\langle g_{h},p_{h}\rangle}{\|g_{h}\|_{{\mathrm{Q}}^{\prime}_{h}}}=\beta_{d}\|p_{h}\|_{\widehat{\mathrm{Q}}(\mathcal{T}_{h})},$
@@
(4.8) $\|(\boldsymbol{u}_{h},q_{h})\|_{\mathbf{V}\times{{\widehat{\mathrm{Q}}(\mathcal{T}_{h})}}}\leq\hat{C}\max\bigl\{\|\boldsymbol{f}\|_{0,\Omega}+\|g\|_{0,\Omega}+\|g\|^{2}_{0,\Omega},(\|\boldsymbol{f}\|_{0,\Omega}+\|g\|_{0,\Omega}+\|g\|^{2}_{0,\Omega})^{2}\bigr\}.$
@@
The discrete well-posedness analysis will use the discrete version of Theorem 3.1, detailed in [12, Theorem 4.1]. The local Lipschitz continuity of the discrete counterpart of $\{A+C\}$ follows similarly as in Theorem 3.4, and the discrete inf-sup condition follows from Lemma 4.4. Hence, to complete the rest of the proof, it is enough to prove the uniformly strong monotonicity.
@@
$=\int_{\Omega}\!\kappa^{-1}|\boldsymbol{u}_{h}-\boldsymbol{v}_{h}|^{2}\!+\mathtt{F}\int_{\Omega}\left((\boldsymbol{u}_{h}+\boldsymbol{w}_{h})|\boldsymbol{u}_{h}+\boldsymbol{w}_{h}|-(\boldsymbol{v}_{h}+\boldsymbol{w}_{h})|\boldsymbol{v}_{h}+\boldsymbol{w}_{h}|\right)\cdot(\boldsymbol{u}_{h}-\boldsymbol{v}_{h}).$
@@
$\mathtt{F}\int_{\Omega}\left((\boldsymbol{u}_{h}+\boldsymbol{w}_{h})|\boldsymbol{u}_{h}+\boldsymbol{w}_{h}|-(\boldsymbol{v}_{h}+\boldsymbol{w}_{h})|\boldsymbol{v}_{h}+\boldsymbol{w}_{h}|\right)\cdot(\boldsymbol{u}_{h}-\boldsymbol{v}_{h})$ $\geq\hat{C}\,\mathtt{F}\|\boldsymbol{u}_{h}-\boldsymbol{v}_{h}\|^{3}_{0,3,\Omega}.$
Finally, setting $\widetilde{\alpha}^{\frac{1}{2}}=\min\{1,\hat{C}^{\frac{1}{3}}\},$ there holds
@@ 4.2 A priori error analysis
In this section, we estimate the error between the exact and approximate solutions, showing that the convergence is robust in the material parameters.
Below, we prove a theorem on a priori error estimate for the Darcy–Forchheimer model, which are robust with respect to $\kappa$, $h$, and $\mathtt{F}$.
@@
where $C_{g}:=\max\{\frac{1}{\widetilde{\alpha}},\frac{1}{\widetilde{\alpha}}(1+\frac{1}{\beta_{d}})((1+\widetilde{\alpha})+\|\boldsymbol{u}\|_{\mathbf{V}}+\|\boldsymbol{u}_{h}^{g}\|_{\mathbf{V}})\}$, which is obtained by using that $\mathrm{dist}(\boldsymbol{u},\mathbf{Z}_{h}^{g})\leq(1+\frac{1}{\beta_{d}})\,\mathrm{dist}(\boldsymbol{u},\mathbf{V}_{h})$ (see for instance, [16, Theorem 2.6]).
@@
$=\sup_{\boldsymbol{0}\neq\boldsymbol{v}_{h}\in\mathbf{V}_{h}}\frac{b_{h}(\boldsymbol{v}_{h},p-p_{h}^{*})+b_{h}(\boldsymbol{v}_{h},p_{h}-p)}{\|\boldsymbol{v}_{h}\|_{\mathbf{V}}}$
$=\sup_{\boldsymbol{0}\neq\boldsymbol{v}_{h}\in\mathbf{V}_{h}}\frac{b_{h}(\boldsymbol{v}_{h},p-p_{h}^{*})+{a(\boldsymbol{u}_{h}-\boldsymbol{u},p-p_{h})}}{\|\boldsymbol{v}_{h}\|_{\mathbf{V}}}$
$\leq\|p-p_{h}^{*}\|_{{\widehat{\mathrm{Q}}(\mathcal{T}_{h})}}+(1+\|\boldsymbol{u}\|_{\mathbf{V}}+\|\boldsymbol{u}_{h}^{g}\|_{\mathbf{V}})\|\boldsymbol{u}-\boldsymbol{u}_{h}\|_{\mathbf{V}},$
which upon using the triangle inequality, shows the following estimate on the pressure
@@
where $\widetilde{C}_{g}:=\max\{(1+\frac{1}{\beta_{d}}+C_{g}^{\frac{1}{2}}(1+\|\boldsymbol{u}\|_{\mathbf{V}}+\|\boldsymbol{u}_{h}^{g}\|_{\mathbf{V}})),C_{g}^{\frac{1}{2}}(1+\|\boldsymbol{u}\|_{\mathbf{V}}+\|\boldsymbol{u}_{h}^{g}\|_{\mathbf{V}})\}$. This concludes the proof.
Additionally, if $\mathbf{Z}_{h}\subseteq\mathbf{Z},$ then, we obtain the following pressure-independent estimate (cf. [16, Theorem B.2]) for the velocity.
@@
where $C_{D}:=\frac{1}{\widetilde{\alpha}}((1+\widetilde{\alpha})+\|\boldsymbol{u}\|_{\mathbf{V}}+\|\boldsymbol{u}_{h}^{g}\|_{\mathbf{V}})(1+\frac{1}{\beta_{d}})>0$ is independent of $h,\kappa,$ and $\mathtt{F}$.
@@
$\widetilde{\alpha}\|\boldsymbol{u}_{h}-\boldsymbol{u}_{h}^{g}\|^{2}_{\mathbf{V}}$
$=a(\boldsymbol{u}_{h}-\boldsymbol{u}_{h}^{g}-\boldsymbol{u}_{h}^{g},\boldsymbol{u}_{h}-\boldsymbol{u}_{h}^{g})+c(\boldsymbol{u}_{h}-\boldsymbol{u}_{h}^{g}+\boldsymbol{u}_{h}^{g};\boldsymbol{u}_{h}-\boldsymbol{u}_{h}^{g}+\boldsymbol{u}_{h}^{g},\boldsymbol{u}_{h}-\boldsymbol{u}_{h}^{g})$
(4.17) $\qquad-a(\boldsymbol{u}_{h}^{g},\boldsymbol{u}_{h}-\boldsymbol{u}_{h}^{g})-c(\boldsymbol{u}_{h}^{g};\boldsymbol{u}_{h}^{g},\boldsymbol{u}_{h}-\boldsymbol{u}_{h}^{g})$
$=a(\boldsymbol{u}_{h},\boldsymbol{u}_{h}-\boldsymbol{u}_{h}^{g})+c(\boldsymbol{u}_{h};\boldsymbol{u}_{h},\boldsymbol{u}_{h}-\boldsymbol{u}_{h}^{g})-a(\boldsymbol{u}_{h}^{g},\boldsymbol{u}_{h}-\boldsymbol{u}_{h}^{g})-c(\boldsymbol{u}_{h}^{g};\boldsymbol{u}_{h}^{g},\boldsymbol{u}_{h}-\boldsymbol{u}_{h}^{g}).$
@@
Recalling from [16, Theorem 2.6] that $\mathrm{dist}(\boldsymbol{u},\mathbf{Z}_{h}^{g})\leq(1+\frac{1}{\beta_{d}})\mathrm{dist}(\boldsymbol{u},\mathbf{V}_{h})$,where $\beta_{d}$ is the discrete inf-sup constant, we achieve the desired pressure-robustness.
Let us remark that by just considering the spaces $\mathbf{H}^{3}_{\Gamma_{\boldsymbol{u}}}({\mathop{\mathrm{div}}\nolimits},\Omega)$ for the velocity and $\mathrm{L}^{2}(\Omega)$ for the pressure and following the approach of [31] one can obtain the pressure and velocity error estimates separately, but they are not necessarily robust with respect to the parameters $\kappa,\mathtt{F}$.
@@
$({\textbf{AP}}^{\boldsymbol{u}}_{h})$. For the Fortin operator $\Pi_{h}:\mathbf{W}^{m,r}(\Omega)\cap\mathbf{H}^{l}(\mathop{\mathrm{div}}\nolimits,\Omega)\to\mathbb{RT}_{k}(\mathcal{T}_{h}),\,0\leq m\leq l,\,1\leq l\leq k+1$ (see [17, 7, 31]), there holds:
$\star\|\Pi_{h}\boldsymbol{v}-\boldsymbol{v}\|_{m,r,\Omega}\leq\acute{C}h^{l-m}\|\boldsymbol{v}\|_{l,r,\Omega}\qquad\forall\boldsymbol{v}\in{\mathbf{W}}^{m,r}(\Omega),\,0\leq m\leq l,\,1\leq l\leq k+1,$
$\star\|\mathop{\mathrm{div}}\nolimits(\Pi_{h}\boldsymbol{v}-\boldsymbol{v})\|_{0,\Omega}\leq\acute{C}h^{l}\|\mathop{\mathrm{div}}\nolimits\boldsymbol{v}\|_{l,\Omega}\qquad\forall\boldsymbol{v}\in\mathbf{H}^{l}(\mathop{\mathrm{div}}\nolimits,\Omega),\,\,0\leq l\leq k+1,$
where $\acute{C}>0$ is independent of $h.$
$({\textbf{AP}}^{p}_{h})$. Let $P_{h}:\mathrm{W}^{l,r}(K)\to\mathbb{P}_{k}(\mathcal{T}_{h}),\,0\leq m\leq l,\,0\leq l\leq k+1$, be the orthogonal $\mathrm{L}^{2}$ projection. Then, there exists a positive constant ${C}_{\sharp}>0,$ independent of $h,$ satisfying (see, [13])
$\star|P_{h}q-q|_{m,r,{K}}\leq{C}_{\sharp}h^{l-m}|p|_{l,r,{K}}\quad\forall q\in\mathrm{W}^{l,r}({K}),{K\in\mathcal{T}_{h}},0\leq m\leq l,\,0\leq l\leq k+1.$
To conclude this section, the following theorem provides the theoretical rate of convergence of the mixed finite element Galerkin scheme (2.3), under suitable regularity assumptions on the exact solution.
@@
Let $(\boldsymbol{u},p)\in\mathbf{V}\times\mathrm{Q}$ be the solution of the continuous problem (2.3) with the following regularity conditions:
$\boldsymbol{u}\in\kappa^{-\frac{1}{2}}\mathbf{H}^{l+1}(\Omega)\cap$ $\mathtt{F}^{\frac{1}{3}}\mathbf{W}^{l+1,3}(\Omega),\quad\qquad\quad\mathop{\mathrm{div}}\nolimits\boldsymbol{u}\in{\mathrm{H}}^{l+1}(\Omega),$
and $p\in\kappa^{\frac{1}{2}}\mathrm{H}^{l+2}(\Omega)\cap\mathtt{F}^{-\frac{1}{3}}\mathrm{W}^{l+2,\frac{3}{2}}(\Omega),$
where $0\leq l\leq k+1$. Then, there exists a positive constant $C_{\flat}>0$, independent of $h,\kappa$, and $\mathtt{F}$, such that
$\|\boldsymbol{u}-\boldsymbol{u}_{h}\|_{\mathbf{V}}+\|p-p_{h}\|_{\widehat{\mathrm{Q}}(\mathcal{T}_{h})}$ $\leq C_{\flat}h^{l+1}\bigl\{\kappa^{-\frac{1}{2}}\|\boldsymbol{u}\|_{l+1,\Omega}+\|{\mathop{\mathrm{div}}\nolimits}\boldsymbol{u}\|_{l+1,\Omega}+\mathtt{F}^{\frac{1}{3}}\|\boldsymbol{u}\|_{l+1,3,\Omega}\bigr\}$
$\ +C_{\flat}h^{l+1}\{|p|^{2}_{l+1,\Omega}+\kappa|p|^{2}_{l+2,\Omega}+\mathtt{F}^{-\frac{2}{3}}|p|_{l+2,\frac{3}{2},\Omega}\}^{\frac{1}{2}}.$

### v2

@@ 4.1 Robust discrete solvability
First, we prove the discrete inf-sup condition, which is an essential ingredient of the proof of the parameter-robust unique solvability Theorem, discussed in this subsection.
@@
For the existence of ${\mathcal{S}_{d}}\in{\mathcal{L}}\bigl((\mathrm{W}_{h,\Gamma_{p}}^{1,q}(\Omega))^{\prime},{\mathbf{L}}^{{q^{\prime}}}_{h}(\Omega)\bigr)$ with $q\in\{2,{r^{\prime}}\}$, let $\phi_{h}\in\mathrm{W}^{1,{q^{\prime}}}_{h,\Gamma_{p}}(\Omega)$ for a given $g_{h}\in(\mathrm{W}_{h,\Gamma_{p}}^{1,q}(\Omega))^{\prime}$ with $\frac{1}{q}+\frac{1}{q^{\prime}}=1$ be a unique solution of the following weak formulation
@@
owing to the Banach–Nečas–Babuška theorem for elliptic problems in Banach spaces [14]. Then, it suffices to define $\mathcal{S}_{d}(g_{h}):=\nabla_{h}\phi_{h}$, where $\nabla_{h}\phi_{h}\in\mathbf{L}_{h}^{{q^{\prime}}}(\Omega)$. The continuous dependence on data coming from (4.8) also implies that $\|\mathcal{S}_{d}(g_{h})\|_{0,q^{\prime};\Omega}\lesssim\|g_{h}\|$. This, in turn, shows $\mathcal{S}_{d}\in{\mathcal{L}}\bigl((\mathrm{W}^{1,q}_{h,\Gamma_{p}}(\Omega))^{\prime},{\mathbf{L}}_{h}^{q^{\prime}}(\Omega)\bigr)$, for $q=2$ and $q=r^{\prime}$ corresponding to $q^{\prime}=2$ and $q^{\prime}=r$, respectively.
@@
$\|q_{h}\|_{{\widehat{\mathrm{Q}}(\mathcal{T}_{h})}}\leq{\beta_{d}}\sup_{\boldsymbol{v}_{h}\in\mathbf{V}_{h}}\frac{|(q_{h},\mathop{\mathrm{div}}\nolimits\boldsymbol{v}_{h})|}{\|\boldsymbol{v}_{h}\|_{\mathbf{V}}}.$
@@
$\sup_{\boldsymbol{0}\neq\boldsymbol{v}_{h}\in\mathbf{V}_{h}}\!\!\frac{|({\mathop{\mathrm{div}}\nolimits}\,\boldsymbol{v}_{h},p_{h})|}{\|\boldsymbol{v}_{h}\|_{\mathbf{V}}}$ $\geq\sup_{g_{h}\in\mathrm{Q}_{h}^{\prime}}\frac{|({\mathop{\mathrm{div}}\nolimits}[\mathcal{S}_{d}(g_{h})],p_{h})|}{\|\mathcal{S}_{d}(g_{h})\|_{\mathbf{V}}}=\sup_{g_{h}\in\mathrm{Q}_{h}^{\prime}}\frac{|\langle g_{h},p_{h}\rangle|}{\|\mathcal{S}_{d}(g_{h})\|_{\mathbf{V}}}$
$\geq\|\mathcal{S}_{d}\|^{-1}_{\mathcal{L}(\mathrm{Q}_{h}^{\prime},\mathbf{V}_{h})}\sup_{g_{h}\in\mathrm{Q}_{h}^{\prime}}\frac{|\langle g_{h},p_{h}\rangle|}{\|g_{h}\|_{{\mathrm{Q}}^{\prime}_{h}}}=\beta_{d}\|p_{h}\|_{\widehat{\mathrm{Q}}(\mathcal{T}_{h})},$
@@
(4.9) $\|(\boldsymbol{u}_{h},q_{h})\|_{\mathbf{V}\times{{\widehat{\mathrm{Q}}(\mathcal{T}_{h})}}}\leq\hat{C}\max\bigl\{\|\boldsymbol{f}\|_{0,\Omega}+\|g\|_{0,\Omega}+\|g\|^{{r-1}}_{0,\Omega},(\|\boldsymbol{f}\|_{0,\Omega}+\|g\|_{0,\Omega}+\|g\|^{{r-1}}_{0,\Omega})^{r-1}\bigr\}.$
@@
The discrete well-posedness analysis uses the discrete version of Theorem 3.1, detailed in [13, Theorem 4.1]. The local Lipschitz continuity of the discrete counterpart of $\{A+C\}$ follows similarly as in Theorem 3.4, and the discrete inf-sup condition follows from Lemma 4.4. Hence, to complete the rest of the proof, it is enough to prove the uniformly strong monotonicity.
@@
$=\int_{\Omega}\!\kappa^{-1}|\boldsymbol{u}_{h}-\boldsymbol{v}_{h}|^{2}$
$\quad+\mathtt{F}\int_{\Omega}\left((\boldsymbol{u}_{h}+\boldsymbol{w}_{h})|\boldsymbol{u}_{h}+\boldsymbol{w}_{h}|^{r-2}-(\boldsymbol{v}_{h}+\boldsymbol{w}_{h})|\boldsymbol{v}_{h}+\boldsymbol{w}_{h}|^{r-2}\right)\cdot(\boldsymbol{u}_{h}-\boldsymbol{v}_{h}).$
@@
$\mathtt{F}\int_{\Omega}\!\!\!\left((\boldsymbol{u}_{h}+\boldsymbol{w}_{h})|\boldsymbol{u}_{h}+\boldsymbol{w}_{h}|^{r-2}\!-\!(\boldsymbol{v}_{h}+\boldsymbol{w}_{h})|\boldsymbol{v}_{h}+\boldsymbol{w}_{h}|^{r-2}\right)\!\cdot(\boldsymbol{u}_{h}\!-\boldsymbol{v}_{h})\gtrsim\mathtt{F}\|\boldsymbol{u}_{h}\!-\boldsymbol{v}_{h}\|^{r}_{0,r,\Omega},$}
and rearranging terms we get
@@ 4.2 A priori error analysis
Now we prove a theorem on a priori error estimates for the Darcy–Forchheimer model, which are robust with respect to $\kappa$, $h$, and $\mathtt{F}$. We recall, for a given subspace $S_{h}$ of the generic Banach space $(S,\|\bullet\|_{S})$, the notation $\mathrm{dist}(v,S_{h}):=\inf_{w\in S_{h}}\|v-w\|_{S}$.
@@
where $C_{g}:={\frac{1}{\widetilde{\alpha}}(1+\frac{1}{\beta_{d}})((1+\widetilde{\alpha})+\|\boldsymbol{u}\|_{\mathbf{V}}+\|\boldsymbol{u}_{h}^{g}\|_{\mathbf{V}})}$, which is obtained by using that $\mathrm{dist}(\boldsymbol{u},\mathbf{Z}_{h}^{g})\leq(1+\frac{1}{\beta_{d}})\,\mathrm{dist}(\boldsymbol{u},\mathbf{V}_{h})$ (see for instance, [17, Theorem 2.6]).
@@
$\quad=\sup_{\boldsymbol{0}\neq\boldsymbol{v}_{h}\in\mathbf{V}_{h}}\frac{b_{h}(\boldsymbol{v}_{h},p-p_{h}^{*})+b_{h}(\boldsymbol{v}_{h},p_{h}-p)}{\|\boldsymbol{v}_{h}\|_{\mathbf{V}}}$
$\quad=\sup_{\boldsymbol{0}\neq\boldsymbol{v}_{h}\in\mathbf{V}_{h}}\frac{b_{h}(\boldsymbol{v}_{h},p-p_{h}^{*})+{a(\boldsymbol{u}_{h}-\boldsymbol{u},\boldsymbol{v}_{h})+c(\boldsymbol{u}_{h};\boldsymbol{u}_{h},\boldsymbol{v}_{h})-c(\boldsymbol{u};\boldsymbol{u},\boldsymbol{v}_{h})}}{\|\boldsymbol{v}_{h}\|_{\mathbf{V}}}$
$\quad\leq\|p-p_{h}^{*}\|_{{\widehat{\mathrm{Q}}(\mathcal{T}_{h})}}+(1+\|\boldsymbol{u}\|_{\mathbf{V}}+\|\boldsymbol{u}_{h}^{g}\|_{\mathbf{V}})\|\boldsymbol{u}-\boldsymbol{u}_{h}\|_{\mathbf{V}},$
where we have used the boundedness of $b_{h}(\bullet,\bullet)$ and $a(\bullet,\bullet)$, the local Lipschitz continuity of the map $\boldsymbol{u}\mapsto|\boldsymbol{u}|^{r-2}\boldsymbol{u}$, and the boundedness of the discrete solution. Then we choose $p_{h}^{*}$ as the best approximation, use triangle inequality, and employ the velocity estimate to obtain
@@
where $\widetilde{C}_{g}:={1+\frac{1}{\beta_{d}}+C_{g}^{\frac{1}{2}}(1+\|\boldsymbol{u}\|_{\mathbf{V}}+\|\boldsymbol{u}_{h}^{g}\|_{\mathbf{V}})}$. This concludes the proof.
Additionally, if $\mathbf{Z}_{h}\subseteq\mathbf{Z},$ then, we obtain the following pressure-independent estimate (cf. [17, Theorem B.2]) for the velocity.
@@
where $0<C_{D}:=\frac{1}{\widetilde{\alpha}}((1+\widetilde{\alpha})+\|\boldsymbol{u}\|_{\mathbf{V}}+\|\boldsymbol{u}_{h}^{g}\|_{\mathbf{V}})(1+\frac{1}{\beta_{d}})<\infty$} is independent of $h,\kappa,$ and $\mathtt{F}$.
@@
(4.18) $\widetilde{\alpha}\|\boldsymbol{u}_{h}-\boldsymbol{u}_{h}^{g}\|^{2}_{\mathbf{V}}$
${\leq}a(\boldsymbol{u}_{h},\boldsymbol{u}_{h}-\boldsymbol{u}_{h}^{g})+c(\boldsymbol{u}_{h};\boldsymbol{u}_{h},\boldsymbol{u}_{h}-\boldsymbol{u}_{h}^{g})-a(\boldsymbol{u}_{h}^{g},\boldsymbol{u}_{h}-\boldsymbol{u}_{h}^{g})-c(\boldsymbol{u}_{h}^{g};\boldsymbol{u}_{h}^{g},\boldsymbol{u}_{h}-\boldsymbol{u}_{h}^{g}).$
@@
Recalling from [17, Theorem 2.6] that $\mathrm{dist}(\boldsymbol{u},\mathbf{Z}_{h}^{g})\leq(1+\frac{1}{\beta_{d}})\mathrm{dist}(\boldsymbol{u},\mathbf{V}_{h})$,where $\beta_{d}$ is the discrete inf-sup constant, we achieve the desired pressure-robustness.
Let us remark that by just considering the spaces $\mathbf{H}^{r}_{\Gamma_{\boldsymbol{u}}}({\mathop{\mathrm{div}}\nolimits},\Omega)$ for the velocity and $\mathrm{L}^{2}(\Omega)$ for the pressure and following the approach of [32] one can obtain the pressure and velocity error estimates separately, but they are not necessarily robust with respect to the parameters $\kappa,\mathtt{F}$.
@@
Let $t\in(1,\infty)$ and denote by $\Pi_{h}:\mathbf{W}^{l,t}(\Omega)\cap\mathbf{H}(\mathop{\mathrm{div}}\nolimits_{1,t},\Omega)\to\mathbb{RT}_{k}(\mathcal{T}_{h})$ the Fortin operator and by $P_{h}:\mathrm{W}^{l,t}(K)\to\mathbb{P}_{k}(K)$, the local orthogonal $\mathrm{L}^{2}$ projection. Then, as a standard consequence of the Bramble–Hilbert lemma, optimal approximation, and inverse inequalities on polynomial spaces, we can assert that
(4.19a) ${\|\boldsymbol{v}-\Pi_{h}\boldsymbol{v}\|_{m,t,\Omega}\leq C_{1}h^{l-m}|\boldsymbol{v}|_{l,t,\Omega}\quad\forall\boldsymbol{v}\in\mathbf{W}^{l,t}(\Omega),\ 1\leq l\leq k+1,}$
(4.19b) ${\|\mathop{\mathrm{div}}\nolimits(\boldsymbol{v}-\Pi_{h}\boldsymbol{v})\|_{m,t,\Omega}\leq C_{2}h^{l-m}|\mathop{\mathrm{div}}\nolimits\boldsymbol{v}|_{l,t,\Omega}\quad\forall\boldsymbol{v}\in\mathbf{W}^{1,t}(\Omega),\ \text{with}}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad{\mathop{\mathrm{div}}\nolimits(\boldsymbol{v})\in\mathrm{W}^{1,t}(\Omega)\ \text{and}\ 0\leq l\leq k+1,}$
(4.19c) ${\|q-P_{h}q\|_{m,t,\Omega}\leq C_{2}h^{l-m}|p|_{l,t,\Omega}\quad\forall q\in\mathrm{W}^{l,t}(\Omega),\ 0\leq l\leq k+1,}$
(4.19d) ${\|\nabla(q-P_{h}q)\|_{0,t,{K}}\leq C_{3}h_{K}^{l}|p|_{l+1,t,K}\quad\forall q\in\mathrm{W}^{l+1,t}({K}),\ K\in\mathcal{T}_{h},\ l\leq k,}$
for $0\leq m\leq l$, where $C_{1},C_{2}>0$ are independent of $h$ and $C_{3}>0$ is independent of $h_{K}$ (see [18, Sections 4.1 and 4.5] for the first three, and [14] for the fourth one).
To conclude this section, the following theorem provides the theoretical rate of convergence for (2.3), under suitable regularity assumptions on the exact solution.
@@
Let $(\boldsymbol{u},p)\in\mathbf{V}\times\mathrm{Q}$ be the solution of the continuous problem (2.3) with the following additional regularity conditions:
$\boldsymbol{u}\in\kappa^{-\frac{1}{2}}{\mathbf{H}^{l}}(\Omega)\cap{\mathtt{F}^{\frac{1}{r}}\mathbf{W}^{l,r}(\Omega)},\quad\mathop{\mathrm{div}}\nolimits\boldsymbol{u}\in{\mathrm{H}^{l}}(\Omega),\quad\text{and}\quad{p=p_{1}+p_{2}+p_{3}},$
$\text{with}\quad{p_{1}\in\mathrm{H}^{l}(\Omega),\quad p_{2}\in\kappa^{\frac{1}{2}}\mathrm{H}^{l+1}(\Omega),\quad\text{and}\quad p_{3}\in\mathtt{F}^{-\frac{1}{r}}\mathrm{W}^{l+1,r^{\prime}}(\Omega)},$
where $1\leq l\leq k+1$. Then, there exist positive constants $C_{\sharp},C_{\flat}>0$, independent of $h,\kappa$, and $\mathtt{F}$, such that
$\|\boldsymbol{u}-\boldsymbol{u}_{h}\|_{\mathbf{V}}+\|p-p_{h}\|_{\widehat{\mathrm{Q}}(\mathcal{T}_{h})}$ $\leq{C_{\sharp}h^{l}\bigl[\kappa^{-\frac{1}{2}}\|\boldsymbol{u}\|_{l,\Omega}+\|\mathop{\mathrm{div}}\nolimits\boldsymbol{u}\|_{l,\Omega}+\mathtt{F}^{\frac{1}{r}}\|\boldsymbol{u}\|_{l,r,\Omega}\bigr]}$
$\quad+{C_{\flat}h^{l}\bigl[|p_{1}|_{l,\Omega}+\kappa^{\frac{1}{2}}|p_{2}|_{l+1,\Omega}+\mathtt{F}^{-\frac{1}{r}}|p_{3}|_{l+1,r^{\prime},\Omega}\bigr]}.$